Let f(x)={3x^2, x lessthan or equal to1;ax+b, x graterthan 1
Find the value of a and b so that f is differentiable at x=1.
1
Expert's answer
2020-05-09T11:42:58-0400

"f(x) = 3 x^2" when "x \leq 1." "\implies f(1)=3"

"f(x)=ax+b" when "x>1 \implies \lim_{x \to 1+}f(x)=a+b"

A differentiable function is also continuous. The function "f(x)" is continuous at "x=1" iff "\lim_{x \to 1-}f(x)=\lim_{x \to 1+}f(x)=f(1)," hence "a+b-3=0."

Now,

"f'(1^-)= \lim_{h \to 0} \frac{f(1-h)-f(1)}{-h} = \lim_{h \to 0} \frac{3(1-h)^2-3}{-h} =\\= \lim_{h \to 0} \frac{3h^2-6h}{-h} = \lim_{h \to 0} 6-3h = 6"

"f'(1^+) = \lim_{h \to 0} \frac{f(1+h)-f(1)}{h} = \lim_{h \to 0} \frac{a(1+h)+b-3}{h} =\\= \lim_{h \to 0} \frac{a+b-3+ah}{h} = a"

The functuon is differentiable iff "f'(1^-)=f'(1^+) \implies a=6".

From "a=6" and "a+b-3=0" it follows that "a=6, b=-3" .